Perform the 4 step process of Hypothesis Test!!

Is the mean cost of jeans for Community College students $80? Using the data below run a hypothesis test.

60 150 40 45 70 80 200 85 40 125
20 150 100 85 100
Answer #1

H0: Null Hypothesis: =80

HA: Alternative Hypothesis: 80

From the given data,the following statistics are calculated:

n = 15

= 1350/15 =90

s = 49.3530

SE = s/

= 49.3530/ = 12.7429

Test statistic is:

t = (90 - 80)/12.7429= 0.7848

Take = 0.05

ndf = 15 - 1 = 14

From Table, critical values of t = 2.1448

Since calculated value of t = 0.7848 is less than thecritical value of t = 2.1448, the difference is not significant. Fail to reject null hypothesis.

Conclusion:
The data support the claim that the mean cost of jeans for Community College students is $80.